- List all the factors of 39.
- Which is larger 9/16 or 4/7?
- Explain in complete sentences and using mathematical vocabulary. Which is larger .3 or 3?
All of the work needs to be shown but getting stuck on coming up with the 3 problem's answers. Please respond when you tutor's get a chance. Thanks!

Answer by gonzo(654)

the factors i can think of are:
1,3,13,39
i can't think of any others.
don't think there are any.
-----
9/16 or 4/7.
by the calculator, 4/7 is larger.
to make the denominators the same, multiply 9/16 * (7/7).
then multiply 4/7 * (16/16).
9/16 becomes 9*7/16*7 becomes 63/(16*7)
4/7 becomes 4*16/16*7 becomes 64/(16*7)
since the denominators are the same, the numerators tell the story.
slight edge to 4/7 confirms what the calculator says.
-----
3 is 10 times larger than .3.
if you divide 3 by 10 you will get .3
if you divide 3 by 100 you will get .03
anything to the right of the decimal place is less than 1.
1 is less than 3.
not sure how else to explain it.
in the decimal system which is what we are using when we add and subtract and multiply or divide, each position in the display of the numer has a meaning which is that the position is multiplied by the power of 10.
anything to the left of the decimal place and the power of 10 is positive.
anything to the right of the decimal place and the power is negative.
10 to a negative power is the same as 1/10 to a positive power.
here's how it looks:


the number is 1111111.11111
abcdefghijklm
the letters are for identification purposes.
h is the decimal point which is the dividing line between whole numbers and fractional parts of a whole number
going from the decimal point to the left, each position is 10 times as large as the position to its right.
g = 1 * 10^0 equals 1 * 1
f = 1 * 10^1 equals 1 * 10
e = 1 * 10^3 equals 1 * 1000
each time you go to the left the number is 10 times larger than the number to its right.
going from decimal point to the right, each position is 10 times smaller than the position to its left.
i = 1 * 10^-1 which is the same as 1 * 1/10^1 which equals 1 * .1 or 1/10
j = 1 * 10^-2 which is the same as 1 * 1/10^2 which equals 1 * .01 or 1/100
k = 1 * 10^-3 which is the same as 1 * 1/10^3 which equals 1 * .001 or 1/1000